a firm has a production function y=X1X2. IF THE MINIMUM COST OF THE PRODUCTION AT W1=W2=2 IS 4 WHAT IS THE LEVEL OF OUTPUT Y....?
MAXI Y=w1x1+w2x2
1
Expert's answer
2013-05-16T09:53:01-0400

In economics, a production function relates physical output of a production process to physical inputs or factors of production. The production function is
one of the key concepts of mainstream neoclassical theories, used to define
marginal product and to distinguish allocative efficiency, the defining focus
of economics. The primary purpose of the production function is to address
allocative efficiency in the use of factor inputs in production and the
resulting distribution of income to those factors, while abstracting away from
the technological problems of achieving technical efficiency, as an engineer or
professional manager might understand it.

If y = x1x2, the level of output is maximized, when y' = 0, x1x2' = x1 + x2 = 0
If W1 = W2 = 2, then maxi Y = 2x1 + 2x2 = 0. So, we shouldn't produce these
goods.